HUBBLE SPACE TELESCOPE DAILY REPORT #1266 PERIOD COVERED: 0000Z Daily Status Report as of 320/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/1 5181 (Doppler Imaging of the Chromospheres and Transition Regions in Cool Stars) The High Resolution Spectrograph (Side One) was used to make observations of the chromosphere and transition region of the star HD37605. This star is the most active of the single early K class dwarf stars. There was a reported carrousel stabilization failure, which resulted in the loss of one observation. A subsequent late observation cause a ripple effect, and several other observations were lost. 1.2 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.3 Completed GHRS/2 5172 (Lyman Alpha Clouds) The GHRS/2 was used to obtain UV Spectra of a pair of quasars in Q0107-025A, separated by little more than an arcsecond. A wavelength calibration observations was also made. A carrousel drive motor voltage out-of-limits was reported at 319/1633Z. No other anomalies were noted. 1.4 Completed WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Significant HSTARs: None 2.4 Operations Notes: Each HST subsystem experienced no significant problems. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-